3 What is the magnitude of the potential difference across the 20-Ω resistor? 10Ω 20 Ω 10Ω 5.0 Ω 10Ω 4. Assume all the lightening in the world represents a constant current of lkA across a 250kV potential. What is the global lightening power? P = IV = 1000A * 250,000V = 250MW The circuit below contains four equivalent 502light bulbs across a voltage source V-100V. 5, 6. a. b. c. Rank the light bulbs from dimmest to brightest (least to most power dissipated). What is the power dissipated by Bulb 3? By what factor does the power of Bulb 1 change
Answer #1

Q3. The equivalent resistance of 3 10 ohm resistors in series is= 10 +10+10=30 ohm

this 30 ohm is in parallel with 20 ohm resistor

Again the equivalent resistance = (30 x 20)/(30 +20) = 12 ohm

this 12 ohm is in series with 5 ohm.. so the equivalent resistance = 12 + 5= 17 ohm

the current in the circuit =11/17 A

potential drop across 5 ohm resistor = current x 5 = (11/17) x 5 = 3.24 V

so potential drop across 20 ohm resistor = 11 - 3.24 = 7.76 V ( as they are in series)

all the best in the course work